# Large Language Models for Scholarly Ontology Generation: An Extensive Analysis in the Engineering Field

### v1

@@ Abstract
Ontologies of research topics are crucial for structuring scientific knowledge, enabling scientists to navigate vast amounts of research, and forming the backbone of intelligent systems such as search engines and recommendation systems. However, manual creation of these ontologies is expensive, slow, and often results in outdated and overly general representations. As a solution, researchers have been investigating ways to automate or semi-automate the process of generating these ontologies. This paper offers a comprehensive analysis of the ability of large language models (LLMs) to identify semantic relationships between different research topics, which is a critical step in the development of such ontologies. To this end, we developed a gold standard based on the IEEE Thesaurus to evaluate the task of identifying four types of relationships between pairs of topics: broader, narrower, same-as, and other. Our study evaluates the performance of seventeen LLMs, which differ in scale, accessibility (open vs. proprietary), and model type (full vs. quantised), while also assessing four zero-shot reasoning strategies. Several models have achieved outstanding results, including Mixtral-8×7B, Dolphin-Mistral-7B, and Claude 3 Sonnet, with F1-scores of 0.847, 0.920, and 0.967, respectively. Furthermore, our findings demonstrate that smaller, quantised models, when optimised through prompt engineering, can deliver performance comparable to much larger proprietary models, while requiring significantly fewer computational resources.
@@
Large Language Models , Zero-Shot Learning , Ontology Generation , Research Topics , Scholarly Knowledge , Scientific Knowledge Graphs1 Introduction
Ontologies of research topics, along with other knowledge organization systems (KOS) such as taxonomies and thesauri, are essential for organising and retrieving information from digital libraries. Indeed, major publishers such as Elsevier, SpringerNature, ACM, IEEE, and PubMed, utilise a range of organization systems like the All Science Journal Classification (ASJC), SpringerNature Taxonomy, ACM Computing Classification System111ACM Computing Classification System - https://dl.acm.org/ccs, the IEEE Theasurus, and Medical Subject Headings, to organise their published content [1, 2, 3] and ensuring efficient scientific dissemination. Furthermore, ontologies of research topics play a vital role in supporting a wide range of intelligent systems to navigate and interpret academic literature [4, 5, 6]. These include search engines [6], conversational agents [7], analytics dashboards [8], and recommender systems [5].
Nevertheless, generating and maintaining ontologies of research topics is both expensive and time-consuming, typically requiring the involvement of several domain experts over an extended period of time [9]. Moreover, the ever-increasing volume of literature, with approximately 2.5 million new scientific papers published annually [10, 11], necessitates more frequent updates to ensure these ontologies accurately reflect the latest research developments. Over the past 25 years, Artificial Intelligence (AI) researchers have concentrated on developing methods for the automated or semi-automated generation of such ontologies, with varying degrees of success [12, 13, 9, 14]. Some notable achievements include the semi-automatic pipeline for generating OpenAlex Topics [15] and the automatic development of the Computer Science Ontology [16], along with a few other initiatives. However, most of those ontologies are still manually created because current systems struggle to generate large-scale, detailed representations of research topics. Consequently, this challenge remains far from being fully addressed.
@@
In this paper, we investigate whether LLMs can effectively address the challenges involved in generating ontologies of research topics. Specifically, we conduct a comprehensive analysis of the ability of modern LLMs to identify semantic relationships between research topics. To this end, we introduce IEEE-Rel-1K, a gold standard derived from the IEEE Thesaurus, designed to evaluate the identification of four types of relationships (broader, narrower, same-as, and other) between 1,000 pairs of topics. We evaluated 17 off-the-shelf LLMs on this dataset, varying in scale, accessibility (open or proprietary), and model type (full or quantised). The models included 3 full open models, 10 open quantised models, and 4 proprietary models. For each model, we employed four distinct zero-shot prompting strategies. The first strategy involved a straightforward prompt asking the model to determine the relationship between two topics, and . The second strategy extended this by applying the same method to both the original pair (, ) and the reversed pair (, ), with the results then being combined. The third strategy utilised a chain-of-thought prompt, encouraging the model to reason through the relationship step by step. The fourth strategy was a two-way version of this chain-of-thought approach, where the process is repeated for both the original and reversed pairs before integrating the outcomes.
Our findings indicate that several models exhibit exceptional zero-shot reasoning capabilities for this task. Proprietary models led the way, with Claude 3 Sonnet achieving an impressive 0.967 F1. However, several open models also delivered highly competitive results, particularly when chain-of-thought prompting techniques were employed. Among the open models, the top performer was Dolphin-Mistral-7B (0.920 F1), a Mistral variant trained on the Dolphin dataset. The Mixtral model also exhibited strong performance (0.847 F1). The findings also indicate that prompting strategies can significantly influence model performance. Different prompting strategies led to dramatic variations in performance, with gains of over 0.354 F1 points in some instances.
@@
We provide the complete codebase for our methodologies, including the gold standard222Gold Standard and code for experiments can be found at https://github.com/ImTanay/LLM-Automatic-Ontology-Generation.
The remainder of this paper is structured as follows. Section 2 provides an overview of the literature. Section 3 defines the task, describes the gold standard, and presents the seventeen LLMs employed in this study. Section 4 describes the experiments and provides details on the implementation setup. Section 5 presents and discusses our results. Section 6 provides an analysis of the types of errors made by the models. Finally, in Section 7 we conclude the paper and provide future directions.
@@ 2 Related Work
In this section, we review the literature, by beginning with a discussion on the ontologies of research areas (Section 2.1) and then explore the state-of-the-art methodologies for generating these ontologies (Section 2.2). In this context, an ontology is a formal representation of knowledge within a specific domain, defining the types of entities that exist and the relationships between them. An ontology of research areas applies this concept to the scholarly domain, systematically organising the various fields of study, their sub-disciplines, and their interconnections [18].
@@ 2.1 Ontologies of Research Areas
One crucial aspect of these ontologies is their coverage. For instance, the Medical Subject Headings (MeSH) is a single-field ontology with over 30K concepts in the field of Medicine [2]. It is maintained by the National Library of Medicine (US), and mostly employed to organised content in the Medline digital library. It receives yearly updates with minor amendments to reflect the latest developments of the field.
The field of Computer Science is represented by the ACM Computing Classification System333The ACM Computing Classification System – http://www.acm.org/publications/class-2012 [3] and the Computer Science Ontology (CSO) [19]. The former contains about 2K research topics and it is maintained by the Association for Computing Machinery (ACM) for its digital library. CSO, on the other hand, offers a more comprehensive view with over 14,000 research concepts. CSO was automatically generated using Klink-2 [9], starting from a corpus of 16 million scientific publications; and is being employed by Springer Nature to enhance the metadata quality of their computer science proceedings, including the Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence [20].
@@
Lastly, OpenAlex Topics is another multi-field taxonomy with 4,800 topics, primarily used within OpenAlex, a free and open catalogue of scholarly articles; and currently at its first release [15].
While most existing ontologies and taxonomies are manually curated, a labour-intensive and costly process, recent automatic and semi-automatic generation methods have led to the creation of CSO and OpenAlex Topics. However, these methods have yet to fully harness the potential of advanced LLMs.
@@ 2.2 Generation of Ontologies of research topics
Previous decades have seen the emergence of several automated solutions for ontology creation, yet these have primarily been applied outside the scholarly domain, explaining why scholarly ontologies often remain manually curated. This section will first explore successful general approaches to automatic ontology generation, followed by a look at the limited applications within the scholarly domain.
@@
Early work focused on semi-automatic methods, combining human expertise with computational techniques. Maedche and Staab [21] introduced a method using linguistic and statistical techniques to extract concepts and relationships from text, but heavily relied on human input. Cimiano et al. [22] advanced this with Text2Onto, integrating techniques from machine learning and natural language processing to automate extraction, though it still required substantial human validation. Fully automatic methods emerged with OntoLearn [23], which used lexical resources and statistical methods to reduce human oversight. The same team [24] later improved it with a graph-based algorithm in OntoLearn Reloaded, enhancing accuracy and robustness in complex domains. The integration of deep learning marked a significant advancement for the field of ontology generation. BERT [25] revolutionised language understanding, enhancing concept extraction and relationship identification. BERTopic [26] leveraged embeddings for coherent topic generation, influencing ontology and taxonomy extraction. Le et al. [27] use hyperbolic embeddings to capture the hierarchical structure of concepts. As recent advancements, Chen et al.[28] employed pre-trained models for developing hierarchical taxonomies, obtaining improved results compared to state-of-the-art solutions.
@@
In the last years, we saw the emergence of several approaches specifically tailored for generating ontologies of research topics. For instance, Shen et al. [29] developed a method to discover concepts and arrange them hierarchically. Their resulting extensive taxonomy was then implemented in Microsoft Academic Graph. Another notable method is Klink-2 [9], which extracts large-scale ontologies of research topics and is currently used to generate CSO. The algorithm starts with a set of keywords and analyses their relationship based on the extent of their co-occurrence. It infers semantic relationships between keyword pairs by utilising a modified subsumption method combined with temporal information to assess hierarchical relationships, and topic similarity strategies to identify keywords referring to the same subject. Finally, the OpenAlex team expanded the ASJC structure found in Scopus by incorporating over 4.5K new topics identified through citation clustering [15]. This involved clustering citation networks to identify coherent and thematically related groups of papers, using LLMs to label these groups, and then connecting the newly created concepts to Scopus’ ASJC.
Despite significant advancements in automating ontology generation, the creation of comprehensive and fine-grained ontologies continues to pose a considerable challenge. We hypothesise, however, that the advanced language understanding capabilities of LLMs could play a crucial role in addressing this gap. This paper represents an initial step towards exploring this potential.
@@ 3.1 Task definition
This paper addresses the challenge of identifying semantic relationships between pairs of research topics, denoted in the following as and . We formalise this task as a single-label, multi-class classification problem, where each topic pair is assigned to one of four distinct categories commonly utilised in research topic ontologies [30]: broader, narrower, same-as, or other.
We define the four categories as follows:
@@
other: and do not relate according to the above categories. E.g., software algorithms and cyclones. Also this relationship is symmetric.

### v2

@@ Abstract
Ontologies of research topics are crucial for structuring scientific knowledge, enabling scientists to navigate vast amounts of research, and forming the backbone of intelligent systems such as search engines and recommendation systems. However, manual creation of these ontologies is expensive, slow, and often results in outdated and overly general representations. As a solution, researchers have been investigating ways to automate or semi-automate the process of generating these ontologies. One of the key challenges in this domain is accurately assessing the semantic relationships between pairs of research topics. This paper presents an analysis of the capabilities of large language models (LLMs) in identifying such relationships, with a specific focus on the field of engineering. To this end, we introduce a novel benchmark based on the IEEE Thesaurus for evaluating the task of identifying three types of semantic relations between pairs of topics: broader, narrower, and same-as. Our study evaluates the performance of seventeen LLMs, which differ in scale, accessibility (open vs. proprietary), and model type (full vs. quantised), while also assessing four zero-shot reasoning strategies. Several models with varying architectures and sizes have achieved excellent results on this task, including Mixtral-8×7B, Dolphin-Mistral-7B, and Claude 3 Sonnet, with F1-scores of 0.847, 0.920, and 0.967, respectively. Furthermore, our findings demonstrate that smaller, quantised models, when optimised through prompt engineering, can achieve strong performance while requiring very limited computational resources.
@@
[…]
1 Introduction
Ontologies of research topics, along with other knowledge organization systems (KOS), such as taxonomies and thesauri, play a vital role in organizing and retrieving information from digital libraries [1]. Major publishers and organizations in this domain, such as Elsevier, Springer Nature, ACM, IEEE, and PubMed, employ a variety of classification systems, including the All Science Journal Classification (ASJC), SpringerNature Taxonomy, ACM Computing Classification System111ACM Computing Classification System - https://dl.acm.org/ccs, the IEEE Theasurus, and Medical Subject Headings. These systems are crucial for enabling efficient information discovery and supporting effective scientific dissemination [2, 3, 4]. Furthermore, ontologies of research topics play a vital role in supporting a wide range of intelligent systems to navigate and interpret academic literature [5, 6, 7]. These include search engines [7], conversational agents [8], analytics dashboards [9], and recommender systems [6].
Nevertheless, generating and maintaining ontologies of research topics is both expensive and time-consuming, typically requiring the involvement of several domain experts over an extended period of time [10]. Moreover, the ever-increasing volume of literature, with approximately 2.5 million new scientific papers published annually [11, 12], demands continuous updates to ensure that these ontologies accurately reflect the latest research developments. Over the past 25 years, Artificial Intelligence (AI) researchers have developed several methods for the automated or semi-automated generation of such ontologies, with varying degrees of success [13, 14, 10, 15]. Some notable achievements include the semi-automatic pipeline for generating OpenAlex Topics [16] and the automatic development of the Computer Science Ontology [17], along with a few other initiatives. However, most of the ontologies in this domain are still manually created because current solutions still struggle to generate large-scale, detailed representations of research topics. Consequently, this challenge remains far from being fully addressed.
@@ 1 Introduction
In this paper, we explore the potential of LLMs to facilitate ontology generation by identifying semantic relationships between research topics, with a specific focus on the field of engineering. This task is a fundamental step in the construction of scientific KOSs, and has traditionally been the most labour-intensive phase, often involving extensive deliberation among domain experts [10]. Specifically, we analyse the capability of several state-of-the-art LLMs to detect three semantic relations frequently adopted in such ontologies: broader, narrower, and same-as. In particular, our experimental analysis focuses on identifying such relations between research topics defined in the IEEE Thesaurus, the most widely adopted ontology in electrical and electronic engineering, which also offers broad coverage of computer science. To support this investigation, we introduce IEEE-Rel-1K, a gold-standard dataset derived from the IEEE Thesaurus, which describes the semantic relationships of 1,000 topic pairs.
We evaluated 17 off-the-shelf LLMs on this dataset, varying in scale, accessibility (open or proprietary), and model type (full or quantised). The models included 3 full open models, 10 open quantised models, and 4 proprietary models. For each model, we employed four distinct zero-shot prompting strategies. The first strategy involved a straightforward prompt asking the model to determine the relationship between two topics, and . The second strategy extended this by applying the same method to both the original pair (, ) and the reversed pair (, ), with the results then being combined. The third strategy utilised a Chain-of-Thought (CoT) prompt, encouraging the model to reason through the relationship step by step. The fourth strategy was a two-way version of this CoT approach, where the process is repeated for both the original and reversed pairs before integrating the outcomes.
Our findings show that several models exhibit strong zero-shot reasoning capabilities for this task. Proprietary models performed best, with Claude 3 Sonnet achieving an impressive F1 score of 0.967. Nonetheless, several open models also delivered highly competitive results, particularly when enhanced with advanced prompting techniques based on Chain-of-Thought (CoT) reasoning. Among the open models, the top performer was Dolphin-Mistral-7B (0.920 F1), a Mistral variant fine-tuned on the Dolphin dataset. The Mixtral model also demonstrated solid performance, reaching an F1 score of 0.847.
Our study produced several notable insights. First, many models exhibit strong zero-shot reasoning capabilities for this task. Second, the performance of LLMs is highly dependent on the prompting strategy employed. In particular, the two-way variant of CoT prompting consistently outperforms alternative approaches. In fact, our results show that prompting strategies can lead to substantial performance gains, with improvements exceeding 0.2 F1 points in several instances. Finally, when the optimal prompting strategy described in this paper is adopted, smaller models, which are more scalable and cost-efficient, can achieve performance comparable to that of the best-performing architectures. These findings point to promising directions for the development of optimised and efficient models for ontology generation.
Ultimately, our objective is to develop a detailed, comprehensive ontology of research topics across all domains. To achieve this, we plan to develop and publish field-specific ontologies using the SKOS standard222SKOS - https://www.w3.org/TR/2009/REC-skos-reference-20090818/, subsequently mapping them together. This paper represents the first step in this ambitious direction.
@@
We provide the complete codebase for our methodologies, including the gold standard333Gold Standard and code for experiments can be found at https://github.com/ImTanay/LLM-Automatic-Ontology-Generation.
The remainder of this paper is structured as follows. Section 2 provides an overview of the literature. Section 3 defines the task, describes the gold standard, and presents the seventeen LLMs employed in this study. Section 4 describes the experiments and provides details on the implementation setup. Section 5 presents and discusses our results. Section 6 provides an analysis of the types of errors made by the models. Section 7 discusses the key findings and the limitations. Finally, in Section 8 we conclude the paper and provide future directions.
@@ 2 Related Work
In this section, we review the relevant literature. We begin with a discussion on knowledge organization systems (Section 2.1), then discuss ontologies of research areas (Section 2.2), and finally explore state-of-the-art methodologies for generating these ontologies (Section 2.3).
2.1 Knowledge Organization Systems
Knowledge Organisation Systems (KOSs) are frameworks for structuring information to facilitate efficient knowledge management and retrieval [19]. These systems, varying in structure and functionality (e.g., hierarchical relationships, synonym control), are commonly categorised into four primary types: term lists, taxonomies, thesauri, and ontologies.
A term list constitutes a linear, non-hierarchical compilation of subject headings or descriptors, facilitating the organisation of document collections [20, 21]. In literature, they are alternatively addressed as pick list [22] or as terminology [23]. These KOSs are distinguished by the absence of explicit relational definitions between subjects. A taxonomy arranges classes within a hierarchical framework, with parent-child relationships [24]. They are organised in tree structures, featuring a root node, which subsequently unfolds into sub-branches. A thesaurus extends the taxonomy structure in which subjects are still organised hierarchically and are further characterised by descriptive attributes, including definitions, related terms, and synonyms [22]. Finally, an ontology is a formal and explicit representation of domain knowledge, classifying entities based on their relevant attributes [25]. In practice, ontologies include concepts, objects, and their relations, and represent the most functionally comprehensive type of KOSs [26]. They offer advanced capabilities, including synonym control, establishment of diverse relationship types, and definition of object properties [19].
In the context of this work, an ontology of research areas applies this concept to the scholarly domain, systematically organising the various fields of study, their sub-disciplines, and their connections.
@@ 2.2 Ontologies of Research Areas
The Medical Subject Headings (MeSH) is a single-field ontology with over 30K concepts in the field of Medicine [3]. It is maintained by the National Library of Medicine (US), and mostly employed to organised content in the Medline digital library. It receives yearly updates with minor amendments to reflect the latest developments in the field.
The field of Computer Science is represented by the ACM Computing Classification System444The ACM Computing Classification System – http://www.acm.org/publications/class-2012 [4] and the Computer Science Ontology (CSO) [27]. The former contains about 2K research topics and it is maintained by the Association for Computing Machinery (ACM) for its digital library. CSO, on the other hand, offers a more comprehensive view with over 14,000 research concepts. CSO was automatically generated using Klink-2 [10], starting from a corpus of 16 million scientific publications; and is being employed by Springer Nature to enhance the metadata quality of their computer science proceedings, including the Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence [28].
@@
OpenAlex Topics is another multi-field taxonomy with 4,800 topics, primarily used within OpenAlex, a free and open catalogue of scholarly articles; and currently at its first release [16].
A comprehensive survey on research topic ontologies is available in [1].
Despite their differences, these systems consistently employ hierarchical structures, using broader/narrower relations, and manage synonyms through ‘same-as’ or related links. Our work strategically focuses on these common relations, allowing us to build a robust framework for creating or enhancing existing ontologies. Furthermore, most existing systems are manually curated, and their manual development is labour-intensive, time-consuming, and costly [1]. Although automatic and semi-automatic generation methods have recently emerged, resulting in the creation of KOS such as CSO and OpenAlex Topics, these approaches have yet to fully exploit the potential of LLMs.
Finally, it is worth noting that several of these ontologies are also integrated in large-scale knowledge graphs [29], that describe the metadata of research publications (e.g., SemOpenAlex [30], AIDA-KG [9]) or specific research entities and concepts (e.g., ORKG [31], CS-KG [32], Nano-publications [33]).
@@ 2.3 Generation of Ontologies of research topics
Previous decades have seen the emergence of several automated solutions for ontology creation, yet these have primarily been applied outside the scholarly domain, explaining why scholarly ontologies often remain manually curated. This section is structured into three parts. First, we examine established methodologies for automatic ontology generation, with a focus on traditional natural language processing (NLP) techniques. Second, we broaden the scope to include the application of large language models (LLMs). Finally, we conclude by reviewing the currently limited applications of these techniques in the scholarly domain.
@@
Early work focused on semi-automatic methods, combining human expertise with computational techniques. Maedche and Staab [34] introduced a method using linguistic and statistical techniques to extract concepts and relationships from text, but heavily relied on human input. Cimiano et al. [35] advanced this with Text2Onto, integrating techniques from machine learning and natural language processing to automate extraction, though it still required substantial human validation. Fully automatic methods emerged with OntoLearn [36], which used lexical resources and statistical methods to reduce human oversight. The same team [37] later improved it with a graph-based algorithm in OntoLearn Reloaded, enhancing accuracy and robustness in complex domains. The integration of deep learning marked a significant advancement for the field of ontology generation. BERT [38] revolutionised language understanding, enhancing concept extraction and relationship identification. BERTopic [39] leveraged embeddings for coherent topic generation, influencing ontology and taxonomy extraction. Le et al. [40] use hyperbolic embeddings to capture the hierarchical structure of concepts. As recent advancements, Chen et al. [41] employed pre-trained models for developing hierarchical taxonomies, obtaining improved results compared to state-of-the-art solutions. Some efforts have also focused on the automatic or semi-automatic evolution of ontologies [42, 43].
This foundation has been significantly advanced by recent research leveraging LLMs for ontology learning. Babaei Giglou et al. [44] introduced LLMs4OL, demonstrating how LLMs can extract concepts, hierarchical relations, and non-taxonomic relations from text, addressing the limitations of previous statistical approaches. However, their methodology suffered from critical limitations, particularly in extracting non-taxonomic relations consistently and maintaining semantic coherence across different domains, which introduces potential bias and hallucination risks. Building upon this foundational work, Saeedizade and Blomqvist [45] conducted a comprehensive examination of LLMs’ capabilities across the entire ontology development lifecycle. Their analysis revealed that while LLMs excel at ontology requirements gathering and conceptualisation, they still require substantial human oversight for formalisation tasks. Critically, they show that models struggles with maintaining precise ontological consistency and their inability to autonomously validate ontological structures without human intervention. Kommineni et al. [46] proposed an innovative framework combining human expertise with LLM capabilities for knowledge graph construction. Their approach demonstrated significant efficiency improvements, reducing expert time requirements while maintaining quality. Nevertheless, the study highlighted consistent challenges, including performance inconsistencies across different knowledge domain contexts and potential semantic drift during automated graph generation. Zhang et al. [47] explored the integration of conversational interfaces into ontology engineering and introduced OntoChat, a framework that enables domain experts to construct ontologies through natural language dialogue. While innovative, the approach revealed significant limitations, particularly in scalability for complex ontologies and the introduction of potential communication ambiguities that could compromise ontological precision. Sun et al. [48] systematically evaluated the potential of LLMs to replace traditional taxonomies. Their comprehensive analysis revealed that LLMs demonstrate outstanding capabilities, but they cannot consistently replace curated taxonomies. Moreover, they found significant variations in performance across domain specialisations and challenges in maintaining long-term consistency.
While LLMs offer clear advantages over traditional ontology generation methods, particularly in concept extraction and the identification of hierarchical relationships, producing comprehensive and detailed ontologies remains an open challenge. However, for this work we hypothesise that carefully designed prompting strategies, combined with LLMs’ advanced language processing capabilities, can yield consistent and domain-specific modelisations.
@@
With regard to generating ontologies in the scholarly domain, research endeavours have explored a variety of (semi-)automated methods. For instance, Shen et al. [49] developed a method to discover research concepts and arrange them hierarchically. Their resulting extensive taxonomy was then implemented in Microsoft Academic Graph. Another notable method is Klink-2 [10], which extracts large-scale ontologies of research topics and is currently used to generate CSO. The algorithm starts with a set of keywords and analyses their relationship based on the extent of their co-occurrence. It infers semantic relationships between keyword pairs by utilising a modified subsumption method combined with temporal information to assess hierarchical relationships, and topic similarity strategies to identify keywords referring to the same subject. Finally, the OpenAlex team expanded the ASJC structure found in Scopus by incorporating over 4.5K new topics identified through citation clustering [16]. This involved clustering citation networks to identify coherent and thematically related groups of papers, using LLMs to label these groups, and then connecting the newly created concepts to Scopus’ ASJC.
In conclusion, LLMs are increasingly being used for ontology generation, as well as a range of research-related tasks such as identifying relevant papers [50, 51], assisting with or even automating the creation of literature reviews [52], and enhancing academic writing and referencing [53, 54]. However, their application to research topic ontologies remains significantly underexplored. In this paper, we address this issue by conducting a thorough evaluation of a broad range of recent LLMs on the task of identifying relationships between scientific topics.
@@ 3.1 Task definition
This paper addresses the challenge of identifying semantic relationships between pairs of research topics, denoted in the following as and . We formalise this task as a single-label, multi-class classification problem, where each topic pair is assigned one of the following categories:
@@
other: and do not relate according to the above categories. E.g., software algorithms and cyclones. This category does not represent a semantic relation; rather, it serves to provide the classifier with a mechanism for labelling negative examples.
We focused our extraction on the broader, narrower, and same-as relations, as these constitute a fundamental set for constructing a Knowledge Organization System (KOS) that effectively represents scientific disciplines [55]. As outlined in Sections 2.1 and 2.2, these relations play a critical role in defining hierarchical structures and managing synonymy.